OpenStudy (anonymous):

What will be the value of a cube + b Cube. If a+b=10 and a square + b square = 58 ?!

OpenStudy (gorv):

a^3+b^3=?? if a+b=10 a^2+b^2=58

OpenStudy (gorv):

a^3+b^3=(a+b)(a^2-ab+b^2)

OpenStudy (gorv):

we need to find ab first

OpenStudy (gorv):

(a+b)^2=a^2+b^2+2ab 10^2=58+2ab 100-58=2ab 42=2ab ab=21

OpenStudy (gorv):

now put your values in a^3+b^3=(a+b)(a^2-ab+b^2)

OpenStudy (anonymous):

thanks !